# Imperceptible Rhythm Backdoor Attacks: Exploring Rhythm Transformation for Embedding Undetectable Vulnerabilities on Speech Recognition

### v1

@@ Abstract
Speech recognition is an essential start ring of human-computer interaction, and recently, deep learning models have achieved excellent success in this task. However, when the model training and private data provider are always separated, some security threats that make deep neural networks (DNNs) abnormal deserve to be researched. In recent years, the typical backdoor attacks have been researched in speech recognition systems. The existing backdoor methods are based on data poisoning. The attacker adds some incorporated changes to benign speech spectrograms or changes the speech components, such as pitch and timbre. As a result, the poisoned data can be detected by human hearing or automatic deep algorithms. To improve the stealthiness of data poisoning, we propose a non-neural and fast algorithm called Random Spectrogram Rhythm Transformation (RSRT) in this paper. The algorithm combines four steps to generate stealthy poisoned utterances. From the perspective of rhythm component transformation, our proposed trigger stretches or squeezes the mel spectrograms and recovers them back to signals. The operation keeps timbre and content unchanged for good stealthiness. Our experiments are conducted on two kinds of speech recognition tasks, including testing the stealthiness of poisoned samples by speaker verification and automatic speech recognition. The results show that our method has excellent effectiveness and stealthiness. The rhythm trigger needs a low poisoning rate and gets a very high attack success rate.
@@ 1 Introduction
However, recent research found that exposing classification datasets to malicious training developers may make the deep neural network (DNN) models vulnerable [2]. In some training procedures ($e.g.$, data collection and preparation and model training), the attackers can manipulate the behavior of speech recognition systems by embedding backdoors to DNN models and cause an extreme security risk. The backdoor adversaries poisoned the model to learn the benign and attacker-specific tasks by implanting the backdoor into the target. The adversaries usually generate poisoned samples and alter their ground truth labels with designed triggers for the poisoned task. For inputs containing no trigger, the victim model behaves normally as its clean parallel model. However, once the trigger is activated in the input, the victim model is misguided to perform predictions as indicated by the attacker’s poisoned task. It is difficult to distinguish the backdoored model from its clean version by simply checking the test accuracy with the test dataset.
@@
Can the backdoor trigger in speech recognition avoid automatic detection and sustain naturalness and speech quality? In this paper, we give a positive answer, and we target rhythm as a trigger from the perspective of the components. The rhythm is mainly related to the speed of each syllable [22]. We propose a non-neural and fast algorithm called Random Spectrogram Rhythm Transformation (RSRT) to generate poisoned samples whose rhythms are transformed. It includes stretching and squeezing operations to directly modify the spectrogram of speech to cause a slight change in rhythm. The poisoned spectrogram is then reconstructed into speech using a neural network vocoder, ensuring the converted speech’s naturalness and intelligibility. Numerous studies have shown that the neural network vocoder exhibits good generalization performance [23, 24] for various modified spectrograms. We mainly focus on the keyword spotting (KWS) task and text-independent speech emotion recognition (TSER) task in our work because the slight change of rhythm does not destroy the content and emotion. Finally, we conducted two evaluation metrics on poisoned samples to verify the consistency of speech components. The experiment results demonstrate that the rhythm trigger gains a high attack success rate with a very low poisoning rate. Our contributions can be summarized as follows:
@@ 2.1 Speech Recognition
2.2 Backdoor Attacks
Backdoor attacks [25, 26] are a series of poisoning methods that aim to bury a backdoor to the model parameters. The attackers usually design specific target labels and guarantee the model will learn the relationship between labels and triggers. Depending on whether training data is poisoned, backdoor attacks can be categorized into poisoning-based attacks and clean-based attacks. Although the poisoning-based attacks enhance the effectiveness of backdoor attacks, the poisoned samples are vulnerable to avoiding detection by defenders [27, 28, 29]. To avoid notice of defender while remaining attacks, Turner et al. [30, 31] firstly explored clean-label backdoor attacks that aim to control the model’s behavior by an adversary without poisoning input samples during inference. From the perspective of attackers, we will explain various attack methods.
2.2.1 Poisoned-Label and Clean-Label Attacks
Poisoned and clean-label attacks usually happen in classification models that learn weight optimization by many pairs of samples and labels. The poisoned-label attackers modify both samples and labels, while the clean-label attackers only modify the samples. Chen et al. blended an attacker-specific Hello Kitty pattern as a trigger with images [32]. Eykholt et al. constructed distributed and spread images as triggers to mask sign images [33]. Sarkar et al. [34] utilized existent facial characteristics as triggers. All these modifications are connected to target labels. Turner et al.[35] first explored a method that leveraged adversarial perturbations or generative models to modify some benign images for clean-label attacks.
2.2.2 Sample-Agnostic and Sample-Specific Attacks
The poisoned images are commonly generated by attacker-specified generators with specific trigger patterns. Sample-agnostic attacks are widespread backdoor attacks in which all poisoned samples share the same trigger. However, this characteristic can be easily exploited for backdoor defense [27, 36]. Nguyen et al. [37] proposed applying visible triggers to be contained in different poisoned samples, which is the first sample-specific backdoor attack method. This method embedded trigger patterns in the edge structure of poisoned images, broke many defense assumptions, and led to good effectiveness on many backdoor defenses. Therefore, sample-specific attacks cause a serious security threat and are worth deeper investigations.
2.2.3 Semantic and Non-Semantic Attacks
The attackers may utilize the semantic information in the images and relate semantic regions with target labels for backdoor attacks, and this method is a semantic attack. Bagdasaryan et al. [38, 39] first explored that certain features are $e.g$, green cars or white birds can be related to attacker-specified target labels. It is noted that the semantic information is ”bird” or ”car.” Accordingly, the victim model will wrongly classify images containing pre-defined semantic information without any digital modification during inference.
2.2.4 All-to-One and All-to-All Attacks
According to the type of target labels, poisoning-label backdoor attacks can be categorized as all-to-one attacks and all-to-all attacks. All-to-one attacks assume that poisoned samples share the same target label, regardless of their original labels. On the contrary, poisoned samples may possess different target labels within all-to-all attacks. For example, Doan et al. [40, 41] applied label shifting function assigning as $LS(y)=(y+1)\mod T$. It noted that $T$ is a controllable integer by attackers, and $y$ is label.
2.2.5 Digital and Physical Attacks
Previous attacks usually occur entirely in digital space, and attackers implement triggers by coding, which belongs to totally digital attacks. Physical attacks are involved in generating poisoned samples in physical space. Chen et al. [32] firstly applied a pair of glasses as a physical trigger for poisoning a face recognition system originating in a camera. Wenger et al. [42, 3] also found that colored dot stickers can be treated as physical triggers. However, Li et al. [43] demonstrated that digital triggers are sensitive to space transformation, such as rotation and shrinkage, which will change the location and form of triggers in poisoned samples. Based on this knowledge, they proposed a transformation-based attack enhancement that keeps the improved attacks effective in the physical world.
@@ 3.2.2 Threat Model
: This paper concentrates on poisoning-based backdoor attacks. And there are some basic principles in this scenario. The adversaries can only modify the open-access training dataset to a poisoned dataset. The victim models will be trained on the poisoned dataset, and the user will deploy the models to the working environment. Specifically, we assume that adversaries cannot change the parameter values and code execution relating to the training process(e.g., loss function, learning schedule, or the resulting model.
@@ 3.2.3 The Goal of Adversary
: The attacker’s goals primarily include stealthiness, effectiveness, and robustness. Stealthiness requires that backdoor attacks have the ability to escape human examination and machine detection. Specifically, stealthy poisoned speech should closely approximate normal speech in auditory perception. Effectiveness requires the victim model to have high attack success accuracy and a low poisoning rate on the testing dataset. It is noted that although some methods achieve very high attack success rates, they often require a concerning proportion of poisoned samples. This configuration may lead to bad stealthiness. Robustness requires that backdoor attacks behave well on simple detection means and remain effective under more difficult settings, such as adaptive defenses and physical-world settings.
@@ 3.2.4 Poisoning-Based Backdoor Attacks Pipeline
We first illustrate backdoor attacks by the notions and their definition of backdoor in Table 1. We denote the classifier $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\theta$ signifies model parameters, $\mathcal{X}\in R^{T,C}$ being the instance space, and $\mathcal{Y}=[1,2,...,K]$ being the label space. The $T,C$ represent the sequence length and channel number. Let $\mathcal{F}_{t}:\mathcal{X}\rightarrow\mathcal{X}$ indicates the attacker-specified trigger function and $\mathcal{F}_{y}:\mathcal{Y}\rightarrow\mathcal{Y}$ indicates label shifting function. Before attacking, the clean training dataset is prepared that is signified as $D=\{(x_{i},y_{i})\}^{N}_{i=1}$, then, the attacker will design the poisoned subset that is conducted by $D_{s}=\{\mathcal{F}_{t}(x_{j}),\mathcal{F}_{y}(y_{j})\}^{M}_{j=1}$ where the replaced subset is $D_{r}=\{(x_{k},y_{k})\}^{M}_{k=1}$. Finally, the poisoned dataset is mixed by $D_{p}=(D-D_{r})\cup D_{s}$. Backdoor attacks request the model to optimize $f_{\theta}$ by following the training objective.
@@ 3.3 Attack via Random Spectrogram Rhythm Transformation
The typical speech triggers borrow methods straightforwardly from image backdoor attacks. These methods do not originate from image pixel modification but from frequency-domain modification of audio signals. For example, PIBA trigger [9] adds a short noise clip to the signal. JingleBack trigger [45] applies pitch shift and distortion to the signal. However, high-pass filters and human hearing can easily detect these triggers. The triggers of modifying speech components retain the naturalness of poisoned utterances without adding any external sounds or making complex distortion. PBSM[17] and VSVC[16] proposed to change the pitch and timbre. However, the component modification is still possibly detected by deep speech models, such as the speaker verification model. To tackle this problem. We turn our attention to another component, rhythm. The rhythm refers to the speed of each speech syllable, representing the pace of the spoken language. We propose to change each syllable speed by a simple spectrogram frame-level algorithm. Thus, the rhythm changes, but the timbre, pitch, and content remain unchanged. The algorithm is used in the attack stage via random spectrogram rhythm transformation (RSRT).
@@ 3.3.2 RSRT Methods
(1) Stretching: Given a spectrogram $X^{D,T}=\{x_{i}|i=1,2,...,T\}$ composed of frames, where D is the number of frequency bins, and T is the number of frames $x_{i}$. It assumes that the stretched spectrogram is $Y_{s}=\{y_{j}|j=1,2,...,T_{s}\}$. Given a scale ratio parameter $\gamma_{s}$ and frame repetition count $\sigma_{s}$, each new frame can be derived by following algorithm 1:
@@
(2) Squeezing: It is worth mentioning that maintaining the continuity of the spectrogram is crucial in our proposed squeezing algorithm. As seen in the stretching algorithm, each frame of the $Y_{s}$ spectrogram maintains continuity. Therefore, when choosing those frames to squeeze, we must maintain the same continuity by avoiding randomly choosing the indexes but following an arithmetic sequence that starts from 0, ends at $T$, and has a common difference of $\phi_{c}$. We believe that using a bilinear downsampling algorithm can maintain the continuity of the speech spectrogram.
@@ 4.1 Experimental Setting for KWS Task
Dataset: We use the Google Speech Commands Dataset v2 [48] as our experimental materials. The dataset contains more than 65,000 audios of 35 categories of keywords that can be used for keyword-spotting devices. In our experiments, 23,682 audios of 10 categories(“yes,” “no,” “up,” “down,” “left,” “right,” “on,” “off,” “stop,” and “go”) are used. We divide the dataset into the training, validation, and test sets in a ratio of 95:5:5, where the validation set belongs to the training set. The poisoned samples only exist in the training set.
Victim models: Our experiments were performed on the following four KWS networks: Resnet-34 [49], Attention-LSTM [50], KWS-ViT [51], EAT-S [52], they behave excellent classification performance on the keyword spotting task.
Baseline and Attack Setup: We compare our attack with the latest speech backdoor attacks. They are listed as follows. (1) backdoor attack with pixel pattern(BadNets) [3], (2) position-independent backdoor attack(PIBA) [9], (3) dual adaptive backdoor attack(DABA) [13], (4) ultrasonic voice as trigger(Ultrasonic) [10], (5) pitch boosting and sound masking(PBSM) [17], and (6) voiceprint selection and voice conversion(VSVC) [16].
@@
Training Setup: We trained all the victim models with the same hyper-parameters. The batch size is 64. The weights are optimized by Adam optimizer with a learning rate of 1e-4 and cross-entropy loss function. We trained 30 epochs to make all models converge.
Table 2: Evaluation of backdoor attack methods on KWS task. Each item shows the $AV/ASR/PN$ in the table.

### v2

@@ Abstract
Speech recognition is an essential start ring of human-computer interaction. Recently, deep learning models have achieved excellent success in this task. However, the model training and private data provider are separated sometimes, potential security threats that make deep neural networks (DNNs) abnormal deserve to be researched. In recent years, the typical threats, backdoor attacks, have been researched in speech recognition systems. The existing backdoor methods are based on data poisoning. The attacker adds some incorporated changes to benign speech spectrograms or changes the speech components, such as pitch and timbre. As a result, the poisoned data can be detected by human hearing or automatic deep algorithms. To improve the stealthiness of data poisoning, we propose a non-neural and fast algorithm called Random Spectrogram Rhythm Transformation (RSRT) in this paper. The algorithm combines four steps to generate stealthy poisoned utterances. From the perspective of rhythm component transformation, our proposed trigger stretches or squeezes the mel spectrograms and recovers them back to signals. The operation keeps timbre and content unchanged for good stealthiness. Our experiments are conducted on two kinds of speech recognition tasks, including testing the stealthiness of poisoned samples by speaker verification and automatic speech recognition. The results show that our method has excellent effectiveness and stealthiness. The rhythm trigger needs a low poisoning rate and gets a very high attack success rate.
@@ 1 Introduction
However, recent research found that exposing classification datasets to malicious training developers may make the deep neural networks(DNNs) vulnerable [2]. In some training procedures ($e.g.$, data collection and preparation and model training), the attackers can manipulate the behavior of speech recognition systems by embedding backdoors to DNN models and cause an extreme security risk. The backdoor adversaries poisoned the model to learn the benign and attacker-specific tasks by implanting the backdoor into the target. The adversaries usually generate poisoned samples and alter their ground truth labels with designed triggers for the poisoned task. For inputs containing no trigger, the victim model behaves normally as its clean parallel model. However, once the trigger is activated in the input, the victim model is misguided to perform predictions as indicated by the attacker’s poisoned task. It is difficult to distinguish the backdoored model from its clean version by simply checking the test accuracy with the test dataset.
@@
Can the backdoor trigger in speech recognition avoid automatic detection and sustain naturalness and speech quality?
In this paper, we give a positive answer, and we target rhythm as a trigger from the perspective of the components. The rhythm is mainly related to the speed of each syllable [22]. We propose a non-neural and fast algorithm called Random Spectrogram Rhythm Transformation (RSRT) to generate poisoned samples whose rhythms are transformed. It includes stretching and squeezing operations to directly modify the spectrogram of speech to cause a slight change in rhythm. The poisoned spectrogram is then reconstructed into speech using a neural network vocoder, ensuring the converted speech’s naturalness and intelligibility. Numerous studies have shown that the neural network vocoder exhibits good generalization performance [23, 24] for various modified spectrograms. We mainly focus on the keyword spotting (KWS) task and text-independent speech emotion recognition (TSER) task in our work because the slight change of rhythm does not destroy the content and emotion. Finally, we conducted two evaluation metrics on poisoned samples to verify the consistency of speech components. The experiment results demonstrate that the rhythm trigger gains a high attack success rate with a very low poisoning rate. Our contributions can be summarized as follows:
@@ 2.1 Speech Recognition
Speech Recognition models $C$ aim to predict the categories from signals or spectrograms of utterances. We assume that speech sample is $X^{D,T}$, where $D$ is number of step vector and $T$ is time steps. The $X$ denotes as spectrograms when $D>1$, or it denotes as signals. The models train parameters to predict more preciously by the following cross-entropy loss objective.
$[p_{o=1},p_{o=2},...,p_{o=M}]=C(x)$ (1)
$L_{ce}=-\sum^{M}_{c=1}\mathbf{y}_{o=c}log(p_{o=c})$ (2)
The $p_{o=c}$ is the probability that the model predicts the sample belongs to class c. After sufficient training, the optimized model will predict the label $y_{p}=\arg\max_{i=1}^{10}p_{o=i}$. Recently, deep neural networks based on residual convolution model[25, 26], Long Short Term Memory model(LSTM)[27], and transformer layers[28] has gained effectiveness in speech recognition.
2.2 Backdoor Attacks in Computer Vision
Backdoor attacks are developed early in computer vision, especially in image and text classification. Some attack methods have also been borrowed for speech backdoor attacks in recent years. We mainly introduce the visible and invisible attacks because they are basic methods. Gu et.al discovered the BadNets[3] and first revealed the backdoor security threat in DNNs. Gu defined the main stages to embed the backdoor to the victim models and perform backdoor attacks: (1) Construct the poisoned training dataset with an attacker-specific trigger function. (2) Train the DNNs with the poisoned training dataset, leading to the hidden backdoor buried in the model’s parameters. (3) Active the trigger when the attacker wanted to mislead the model’s prediction in the inference stage. It is noted that the triggers are usually bound to inputted samples. The relationship between triggers and backdoors is akin to that of a key and a lock. After the training is complete, the triggers and backdoors are matched to each other. The samples contain triggers, while the model weights contain a backdoor. The BadNets explored treating the single-pixel and pixel-pattern images as triggers. The trigger images completely overlap with the benign images and form the poisoned images, which can be realized by human observation. In a similar vein, the reflection image[29], a fixedly blended image[30], one malicious pixel[31], and fixed and sinusoidal pinstripes [32]can also be triggers for visible attacks. However, the visible triggers have risks of detection. To satisfy the invisibility requirement, Turner et al.[4] proposed to perturb the amplitude of the benign pixel values with a backdoor trigger instead of by replacing the corresponding pixels with the chosen pattern. The agitation made it difficult to identify the poisoned images. Cheng et al.[33] proposed utilizing style transfer to conduct the invisible attack. Guo et.al.[34] made the attacks invisible by hidden feature triggers. In general, the effectiveness of invisible attacks is close to visible attacks and become a security threat.
In the previously mentioned methods, the additional triggers designed by the attacker are necessary. Lin et.al. [35] proposed directly using the combinations of existing benign subjects or features of training images themselves as the trigger. Since these features represent semantic information, this type of attack can be classified as a semantic backdoor attack[36, 37, 38, 39, 40]. Semantic triggers simply recombine existing semantic features within the image without introducing new noise or images. Therefore, they can resist most defense methods based on trigger elimination[41, 42, 43].
2.3 Backdoor Attacks in Speech Recognition
Typically, executing an effective backdoor attack requires the attacker to be familiar with the data properties of the samples and to design suitable triggers accordingly. The properties of speech extremely differ from images. Image data is typically represented as a three-dimensional pixel matrix with spatial correlations, where neighboring pixels have a certain level of association and image features usually behave in local space. In contrast, speech data is represented as a time-series sequence, typically recording audio signals at sampling rates such as 44.1kHz or 16kHz. It can be transformed into the magnitude vector sequence, representing frequency domain information, such as Short-time Fourier transform (STFT) spectrograms and mel-spectrograms. Considering the characteristics of speech, speech backdoor attacks can be classified into (1) methods based on the addition of extra noisy speech and perturbation on signals(Noise trigger and Perturbation trigger)[44, 10, 8, 9, 45, 13, 46, 14]. (2) methods based on the modification of speech components/elements(Element trigger)[15, 16, 17, 18]. Koffas et.al [44] proposed a series of perturbation operations(e.g., pitch shift, reverberation and chorus) to perform digital music effects as a pertubation trigger. They also utilize ultrasonic sounds[10], which are sharp as the noise trigger. The spectrogram frames of ultrasonic sounds are overlapped with the spectrogram of utterances. Zhai et.al[8] adopted a low-volume one-hot-spectrum noise with different frequencies as noise trigger patterns for speaker verification. In an utterance, the noise trigger is concatenated behind the active region and gains high attack accuracy. However, these triggers have some drawbacks. Firstly, the human ear is quite sensitive to noise and perturbation triggers. Thus, they can be detected by human checking. Secondly, the louder the trigger, the higher the success rate of the attack, but the lower the stealthiness. To tackle these problems, from the perspective that speech is composed of elements such as content, timbre, and fundamental frequency[22, 19], Ye et.al[15, 16] proposed VSVC to treat the timbre as speech backdoor attack trigger. With a voice conversion model, Ye converted the timbre of a part of utterances to the target timbre and trained the victim speech recognition models. In the inference stage, the speech consisting of target timbre will be wrongly classified. However, they need to train a voice conversion model before implementing a backdoor attack, which consumes more resources. Further, Cai et.al[17] proposed PBSM to use the pitch as a trigger. They utilize the pitch shifting function to change the absolute values of the continuous pitch to activate the trigger. Cai et.al[18] also demonstrated that the pitch and timbre triggers could be combined as element triggers for multi-target attacks, which gained excellent attack effectiveness on speech recognition. The element triggers also attached high stealthiness because the elements vary greatly in a recognition dataset. For example, a KWS dataset can be recorded by many speakers.
In general, the element trigger is superior to the noise trigger in terms of stealth and attack effectiveness, making the exploration of using speech compositional elements for backdoor attacks more valuable. It is concerning that modifications to the pitch and timbre could also be detected by fundamental frequency(F0) analysis neural networks[47] and speaker verification systems[48]. Thus, we try to explore the speech element owning better stealthiness.
@@ 3.2.2 Threat Model
This paper concentrates on poisoning-based backdoor attacks. And there are some basic principles in this scenario. The adversaries can only modify the open-access training dataset to a poisoned dataset. The victim models will be trained on the poisoned dataset, and the user will deploy the models to the working environment. Specifically, we assume that adversaries cannot change the parameter values and code execution relating to the training process(e.g., loss function, learning schedule, or the resulting model.
@@ 3.2.3 The Goal of Adversary
The attacker’s goals primarily include stealthiness, effectiveness, and robustness. Stealthiness requires that backdoor attacks have the ability to escape human examination and machine detection. Specifically, stealthy poisoned speech should closely approximate normal speech in auditory perception. Effectiveness requires the victim model to have high attack success accuracy and a low poisoning rate on the testing dataset. It is noted that although some methods achieve very high attack success rates, they often require a concerning proportion of poisoned samples. This configuration may lead to bad stealthiness. Robustness requires that backdoor attacks behave well on simple detection means and remain effective under more difficult settings, such as adaptive defenses and physical-world settings.
@@ 3.2.4 Poisoning-Based Backdoor Attacks Pipeline
We first illustrate backdoor attacks by the notions and their definition of backdoor in Table 1. We denote the classifier $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\theta$ signifies model parameters, $\mathcal{X}\in R^{T,C}$ being the instance space, and $\mathcal{Y}=[1,2,...,K]$ being the label space. The $T,C$ represent the sequence length and channel number. Let $\mathcal{F}_{t}:\mathcal{X}\rightarrow\mathcal{X}$ indicate the attacker-specified trigger function and $\mathcal{F}_{y}:\mathcal{Y}\rightarrow\mathcal{Y}$ indicates label shifting function. Before attacking, the clean training dataset is prepared that is signified as $D=\{(x_{i},y_{i})\}^{N}_{i=1}$, then, the attacker will design the poisoned subset that is conducted by $D_{s}=\{\mathcal{F}_{t}(x_{j}),\mathcal{F}_{y}(y_{j})\}^{M}_{j=1}$ where the replaced subset is $D_{r}=\{(x_{k},y_{k})\}^{M}_{k=1}$. Finally, the poisoned dataset is mixed by $D_{p}=(D-D_{r})\cup D_{s}$. Backdoor attacks request the model to optimize $f_{\theta}$ by following the training objective.
@@ 3.3 Attack via Random Spectrogram Rhythm Transformation
The typical speech triggers borrow methods straightforwardly from image backdoor attacks. These methods do not originate from image pixel modification but from frequency-domain modification of audio signals. For example, PIBA trigger [9] adds a short noise clip to the signal. JingleBack trigger [44] applies pitch shift and distortion to the signal. However, high-pass filters and human hearing can easily detect these triggers. The triggers of modifying speech components retain the naturalness of poisoned utterances without adding any external sounds or making complex distortion. PBSM[17] and VSVC[16] proposed to change the pitch and timbre. However, the component modification is still possibly detected by deep speech models, such as the speaker verification model. To tackle this problem, we turn our attention to another component, rhythm. The rhythm refers to the speed of each speech syllable, representing the pace of the spoken language. We propose to change each syllable speed by a simple spectrogram frame-level algorithm. Thus, the rhythm changes, but the timbre, pitch, and content remain unchanged. The algorithm is used in the attack stage via random spectrogram rhythm transformation (RSRT).
@@ 3.3.2 RSRT Methods
Stretching: Given a spectrogram $X^{D,T}=\{x_{i}|i=1,2,...,T\}$ composed of frames, where D is the number of frequency bins, and T is the number of frames $x_{i}$. It assumes that the stretched spectrogram is $Y_{s}=\{y_{j}|j=1,2,...,T_{s}\}$. Given a scale ratio parameter $\gamma_{s}$ and frame repetition count $\sigma_{s}$, each new frame can be derived by following algorithm 1:
@@
Squeezing: It is worth mentioning that maintaining the continuity of the spectrogram is crucial in our proposed squeezing algorithm. As seen in the stretching algorithm, each frame of the $Y_{s}$ spectrogram maintains continuity. Therefore, when choosing those frames to squeeze, we must maintain the same continuity by avoiding randomly choosing the indexes but following an arithmetic sequence that starts from 0, ends at $T$, and has a common difference of $\phi_{c}$. We believe that using a bilinear downsampling algorithm can maintain the continuity of the speech spectrogram.
@@ 4.1 Experimental Setting for KWS Task
Dataset. We evaluate our method on Google Speech Commands versions 1(GSCv1) and 2 (GSCv2) dataset[52]. Version 1 contains 64,727 utterances from 1,881 speakers for 30 word categories and version 2 has 105,829 utterances from 2,618 speakers for 35 words. Each utterance is 1 sec long, and the sampling rate is 16 kHz. The datasets are pre-processed following[52] for the keyword spotting task where only 10 words are interesting targets, specifically: “Yes”, “No”, “Up”, “Down”, “Left”, “Right”, “On”, “Off”, “Stop” and ”Go”. We divide the dataset into the training, validation, and test sets in a ratio of 95:5:5, where the validation set belongs to the training set. The poisoned samples only exist in the training set. The audio segments are processed by extracting mel-spectrograms, where the window length is 1024, hop length is 256, the FFT bin is 1024, the mel bin is 80.
Victim models. Our experiments were performed on the following four KWS networks: Resnet-34 [53], Attention-LSTM [54], KWS-ViT [55], EAT-S [56], they behave excellent classification performance on the keyword spotting task.
Baseline and Attack Setup. We compare our attack with the latest speech backdoor attacks. They are listed as follows. (1) backdoor attack with pixel pattern(BadNets) [3], (2) position-independent backdoor attack(PIBA) [9], (3) dual adaptive backdoor attack(DABA) [13], (4) ultrasonic voice as trigger(Ultrasonic) [10], (5) pitch boosting and sound masking(PBSM) [17], and (6) voiceprint selection and voice conversion(VSVC) [16].
@@
Training Setup. We trained all the victim models with the same hyper-parameters. The batch size is 64. The weights are optimized by Adam optimizer with a learning rate of 1e-4 and cross-entropy loss function. We trained 30 epochs to make all models converge.
4.2 Experimental Setting for TSER Task
Dataset. We used two speech emotion datasets: Emotional Speech Dataset (ESD) [57] and Interactive emotional dyadic motion capture database (IEMOCAP)[58] for the TSER task. The ESD dataset consists of 350 parallel utterances spoken by 10 native English and 10 native Chinese speakers and covers 5 emotion categories (neutral, happy, angry, sad, and surprise). We only use ESD samples from the English language for training. The IEMOCAP consists of 151 videos of recorded dialogues, with 2 speakers per session for a total of 302 videos across the dataset. Each segment is annotated for the presence of 9 emotions. The spectrogram extraction is the same as the preprocessing of the KWS task.
Victim models. Our experiments were performed on signal process deep neural models. We choose AST [28], SER-AC [59] and SER-CNN [60]. These models use only signal information to learn emotional classification and reach effective performance.
Baseline and Attack Setup. We set the same proposed attacking configuration as the KWS task.
Training Setup. We trained all the victim models with AST’s hyperparameters. The batch size is 12. The weights are optimized by Adam optimizer with a learning rate of 1e-5 and cross-entropy loss function. The learning rate is cut into half every epoch after the 2nd epoch. We trained 30 epochs to make all models converge.
Table 2: Attack results on GSC v1 dataset towards KWS task. Each item shows the $AV/ASR/PN$ in the table.